OpenStudy (rva.raghav):

HW11P1 Qn.2, not getting...If we tune to a station at f=860.0kHz what is the capacitance, in picoFarads, of the tuning capacitor? Qn4.If next we tune to a station at f=1300.0kHz what is the capacitance, in picoFarads, of the tuning capacitor?

OpenStudy (anonymous):

H11P1 1.ω=2πf Capacitance C=1/((ω^2)L) 2.Bandwidth Δω = RL/L + 1/(Rc *C) 3.Capacitance C=1/((ω^2)L) 4.Bandwidth Δω = RL/L + 1/(R*C *C)

OpenStudy (anonymous):

Pls click best response if u get d answer

OpenStudy (anonymous):

with f= 860khz RL=4ohm Rc=490k L=0.65mHenry I can't get the bandwidth to check I get 44.1 with those values, still won't check, :c

OpenStudy (benzo):

Δω = RL/L + 1/(Rc *C) #2.....4.0/0.65 + 1/(490k *44.1) .....=6.15384620012 is no good for me....help please

OpenStudy (anonymous):

check the value of 'C' again.. did u forget to put the value of C's unit..

OpenStudy (rva.raghav):

akash can u help me

OpenStudy (anonymous):

Jone133, try this values: For 860kHz C = 52.6902190593 pF band =7.14386096211 kHz

OpenStudy (anonymous):

For 860kHz C = 52.6902190593 pF band =7.14386096211 kHz For 1300 C = 23.0589858085 pF band = 15.0652906669 kHz If works, please mark as best response :D

OpenStudy (anonymous):

for 770 Khz??

OpenStudy (anonymous):

allwynj, try 5.92114027851

OpenStudy (anonymous):

5.92